### v3

# Dynamical Transition due to Feedback-induced Skin Effect

Abstract
The traditional dynamical phase transition refers to the appearance of singularities in an observable with respect to a control parameter for a late-time state or singularities in the rate function of the Loschmidt echo with respect to time. Here, we study the many-body dynamics in a continuously monitored free fermion system with conditional feedback under open boundary conditions. We surprisingly find a novel dynamical transition from a logarithmic scaling of the entanglement entropy to an area-law scaling as time evolves. The transition, which is noticeably different from the conventional dynamical phase transition, arises from the competition between the bulk dynamics and boundary skin effects. In addition, we find that while quasidisorder or disorder cannot drive a transition for the steady state, a transition occurs for the maximum entanglement entropy during the time evolution, which agrees well with the entanglement transition for the steady state of the dynamics under periodic boundary conditions.
Non-Hermitian physics has sparked considerable interests from various fields in recent years due to the presence of intriguing phenomena [1, 2, 3, 4], such as exceptional points or rings [5, 6, 7, 8, 9] and non-Hermitian skin effects [10, 11, 12, 13, 14, 15, 16]. In fact, non-Hermiticity is ubiquitous in the dynamics of many-body quantum systems [1]. For instance, when considering postselection in continuous measurements, the many-body dynamics is dictated by a non-Hermitian Hamiltonian [1]. Through non-Hermitian many-body dynamics, many interesting non-equilibrium phenomena have been discovered [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 29, 31, 28, 30, 32, 33]. In particular, it has been found that the steady state of a non-Hermitian dynamics of free fermions in the Hatano-Nelson (HN) model is the skin state where particles mainly reside at a half of a system, possessing an area-law scaling of the entanglement entropy [27]. Later, it was shown that disorder or quasidisorder can drive an entanglement phase transition for the steady state with peculiar scaling properties [34, 35, 36].
However, to observe these interesting phenomena, one has to apply postselection, which is formidably challenging for a large system. Fortunately, non-Hermitian dynamics arises naturally in quantum dissipative systems or continuously monitored systems, as dictated by the Lindblad master equation, due to the typically non-Hermitian nature of the Liouvillian that governs the system’s dynamics. Interestingly, the Liouvillian eigenvectors can also exhibit skin effects [37, 38, 39]. In addition, Wang et al. studied the dynamics of a quantum system under continuous measurements with conditional feedback [2] and found that the late-time steady state always concentrates at an edge [2, 3].
In the field of non-equilibrium many-body dynamics, one central question to ask is whether there exists a dynamical phase transition. Such a transition is classified into two types. One type corresponds to the appearance of a non-analytical behavior in an observable with respect to a control parameter for a late-time state [42, 43, 44, 45, 46, 47]. Roughly speaking, measurement induced entanglement phase transitions [48, 49, 50, 52, 53, 54, 55, 56, 57, 51, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] may fall into this category when the entanglement entropy is regarded as an observable and the measurement rate as a control parameter. The other type of dynamical phase transition occurs if the rate function of the return probability develops a singularity as time evolves [68, 69, 70, 71, 72]. Interestingly, it has been shown that the presence of skin effects precludes the existence of the entanglement phase transition for the steady state [2, 3]. We therefore ask whether the many-body skin effects can lead to the second type dynamical phase transition or other novel dynamical phase transition beyond this traditional one with respect to time.
Figure 1: (a) The time evolution of the trajectory averaged bipartite entanglement entropy $S_{L/2}$ for different system sizes $L$ ($t$ denotes the evolution time). The black dashed line highlights the transition point between the logarithmic-law and area-law regimes. The inset shows the data for several $t/L$ values in a linear-log scale with $L$ up to 448, illustrating the log-law scaling. (b) The time evolution of the trajectory averaged density distribution with respect to site $l$ for a system with $L=128$. Inset: the average velocity ${v}(t)$ under OBCs as a function of $t/L$ (red line). The black dashed line represents a linear fit of ${v}(t)$ before the transition point.
Here we study the many-body dynamics of a continuously monitored free fermion system under open boundary conditions (OBCs). Surprisingly, we find the existence of a novel dynamical transition from a logarithmic scaling of the entanglement entropy to an area-law scaling with respect to $t/L$ where $t$ is the evolution time and $L$ is the system size (see Fig. 1), although the late-time states approach the skin state even in the presence of strong quasidisorder [see Fig. 3(d)]. The dynamical transition arises due to the competition between the bulk dynamics and the dynamics significantly affected by edges. The competition can also be seen from the fact that this dynamical transition does not happen under periodic boundary conditions (PBCs). It is important to note that this transition is fundamentally different from the traditional two types of dynamical phase transitions. In our case, instead of the return probability, the probability for the evolving state to be found in an ideal many-body skin state exhibits a sharp rise across the transition point. In addition, we investigate whether quasidisorder can induce an entanglement transition under OBCs in this system. We find that for the late-time steady state, the transition does not happen. However, we show that the dynamics at earlier times is mainly governed by the bulk, resulting in a rapid growth of the entanglement entropy. Later, the entanglement entropy reaches a maximum followed by a decline to a stable value due to the strong boundary effects caused by skin effects. Remarkably, we find that these maximum entanglements undergo a transition from a logarithmic to an area-law scaling as we raise the quasidisorder strength. The results agree well with the transition observed for the steady states under PBCs.
Trajectory evolution.— We start by considering a spinless fermion chain of length $L$ with nearest-neighbor hopping and an onsite quasiperiodic potential described by the Hamiltonian,
$\hat{H}=\sum_{l}\left(J(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+\text{H.c.})+W\cos{(2\pi\alpha l)}\hat{c}_{l}^{\dagger}\hat{c}_{l}\right),$ (1)
where $\hat{c}_{l}^{\dagger}\ \left(\hat{c}_{l}\right)$ is the creation (annihilation) operator of a fermion at site $l$, $J$ denotes the hopping strength (we set $J=1$ as units of energy hereafter), and $W$ represents the amplitude of the potential, which is quasiperiodic when $\alpha$ takes irrational values (we consider $\alpha=\left(\sqrt{5}-1\right)/2$ without loss of generality). To generate skin effects, we follow Refs. [2, 3] to introduce continuous measurements with feedback represented by the quantum jump operator $\hat{L}_{l}=e^{i\theta\hat{n}_{l+1}}\hat{d}_{l}^{\dagger}\hat{d}_{l}$ with the strength proportional to $\gamma$, where $\hat{n}_{l}=\hat{c}_{l}^{\dagger}\hat{c}_{l}$, $\hat{d}_{l}=\frac{1}{\sqrt{2}}(\hat{c}_{l}+i\hat{c}_{l+1})$, and $\theta$ is a parameter that takes real values (we set $\theta=\pi$ hereafter). The evolution trajectory of a pure state $|\psi_{t}\rangle$ at time $t$ is then described by a stochastic Schrödinger equation (SSE) [1, 73],
$d|\psi_{t}\rangle=-i\hat{H}_{\mathrm{eff}}|\psi_{t}\rangle dt+\sum_{l}\left({\hat{L}_{l}}/\sqrt{a_{0}}-1\right)|\psi_{t}\rangle dW_{l},$ (2)
where $\hat{H}_{\mathrm{eff}}=\hat{H}-i\frac{\gamma}{2}\sum_{l}\hat{L}^{\dagger}_{l}\hat{L}_{l}$ is the effective Hamiltonian, $a_{0}=\langle\psi_{t}|\hat{L}^{\dagger}_{l}\hat{L}_{l}|\psi_{t}\rangle$, and $\left\{dW_{l}\right\}$ denotes a set of Poisson random numbers that take the values of either 0 or 1 with the mean value of $\overline{dW_{l}}=\gamma a_{0}dt$. Without loss of generality, we take $\gamma=0.5$ for all subsequent calculations. Previous studies have shown that when $\theta\neq 0$ (e.g., $\theta=\pi$), the steady state is always the skin state for a system under OBCs since the detected right-moving particles will be converted to the left-moving ones through the imposed phase [2, 3]. In contrast to these studies, we will examine whether a transition appears as time evolves.
To study the dynamical transition, we initialize the system in the Neél state, that is, $|\psi_{0}\rangle=\prod_{l=1}^{L/2}\hat{c}_{2l}^{\dagger}|0\rangle$, where $|0\rangle$ is the vacuum state. We then numerically calculate each trajectory of $|\psi_{t}\rangle$ by solving Eq. (2). The observables are calculated in each trajectory at every time slice, and then averaged over all trajectories. Given that the effective Hamiltonian is quadratic and the jump operators have a simple form, the dynamics could be efficiently simulated with a Slater determinant state [73]. Furthermore, the Gaussian structure of the states allows for the extraction of observables from the correlation matrix $C_{ij}(t)=\langle\psi_{t}|\hat{c}^{\dagger}_{i}c_{j}|\psi_{t}\rangle$. To capture the entanglement transition, we focus on the von Neumann entanglement entropy calculated by [74], $S_{A}=-[\textrm{Tr}\left(C_{A}\log{C_{A}}+(1-C_{A})\log{(1-C_{A})}\right)]$, where $C_{A,mn}=\langle\hat{c}_{m}^{\dagger}\hat{c}_{n}\rangle$ with $m,n\in A$ represents the correlation matrix of a subsystem $A$ of an evolving state. Hereafter, we use []delimited-[][$\dots$][ … ] to denote the average over trajectories (our numerical results are averaged over 180-2000 trajectories unless stated otherwise) and use $S_{L/2}$ to denote the entanglement entropy of the left bipartition of the chain.
Figure 2: The time evolution of (a) the probability $f_{\mathrm{skin}}$ of the evolving state $|\psi_{t}\rangle$ being in an ideal many-body skin state $|\psi_{\mathrm{skin}}\rangle$ and (b) the return probability $f_{r}$ for different system sizes averaged over trajectories. In (b), we consider 96 trajectories. The black dashed line marks the transition point estimated in Fig. 1(a). (c) The density distribution with respect to the momentum $k$ for the steady state of the dynamics under PBCs with $L=128$.
Dynamical transition.— We now investigate the entanglement transition that occurs during dynamics for a system under OBCs without quasidisorder (i.e., $W=0$). To compare the dynamics of systems with different sizes $L$, we rescale the time $t$ to $t/L$. Remarkably, we find that the trajectory averaged entanglement entropy $S_{L/2}$ undergoes a transition from a log-law to an area-law regime at $t_{c}/L\approx 0.79$ as shown in Fig. 1(a). Specifically, as time evolves, $S_{L/2}$ first increases, then decreases, and eventually reaches a steady value, which is independent of system sizes. The log-law scaling before the transition is shown in the inset of Fig. 1(a).
To explain the behavior of $S_{L/2}$, we examine the evolution of the density distribution $[\langle\hat{n}_{l}\rangle]$. As shown in Fig. 1(b), all particles move to the left side of the chain and eventually populate the left part. This phenomenon is a result of the feedback-induced skin effect, which has been discussed in Refs. [2, 3]. However, the dynamical transition has not been found there. By a comparison with the evolution of $S_{L/2}$, we observe that before all particles reach the left side, $S_{L/2}$ increases with increasing $L$. Once the state $|\psi_{t}\rangle$ becomes a skin state, $S_{L/2}$ no longer grows.
Figure 3: (a) The phase diagram of the trajectory averaged bipartite entanglement entropy in the plane $(t/L,W)$ with $L=320$. The black line represents the dynamical transition line, separating the log-law regime and the area-law regime. (b) The trajectory averaged entanglement entropy $S_{L/2}$ versus $t/L$ at $W=1$ with $L=64,128,192,256,320,384,448$ (from light to dark colors). Inset: the linear-log scale plot of $S_{L/2}$ versus $L$ at $t/L=0.2,0.5,0.8,1.1$ with a log-law fit (the dashed lines). (c) The correlation functions $C(l)$ versus $(L/\pi)\sin{(\pi l/L)}$ at different $(t/L,W)$ marked by solid blue, red and green circles in (a), respectively. Here $L=320$. The inset displays a data collapse for different system sizes $L=64,128,192,256,320$, with the same axes range as the main plot. (d) The steady-state density distribution ${n}(l)$ at $W=0,3,6$. The total evolution time is set to 10000 to ensure that the final state reaches the steady state.
We note that this dynamical transition is different from the conventional dynamical phase transition that exhibits singularities in the rate function of the Loschmidt echo at the transition point [70]. In our dynamics, the return probability $f_{r}=[|\langle\psi_{0}|\psi_{t}\rangle|^{2}]$ suddenly drops to zero at a very short time as shown in Fig. 2(b). Interestingly, we find that the probability $f_{\mathrm{skin}}=[|\langle\psi_{\mathrm{skin}}|\psi_{t}\rangle|^{2}]$ of the evolving state being in an ideal many-body skin state $|\psi_{\mathrm{skin}}\rangle=\prod_{l=1}^{L/2}\hat{c}_{l}^{\dagger}|0\rangle$ experiences a sudden rise at the transition point [Fig. 2(a)], indicating that the skin state develops through a dynamical transition.
We expect that the dynamical transition arises from the competition between the dynamics under PBCs and boundary skin effects. At earlier times, since most particles are far away from a boundary, their time evolution is mainly governed by the system under PBCs. However, as time progresses, particles approach a boundary so that the boundary effect becomes relevant.
To demonstrate the competition, we analyze the average velocity per particle ${v}(t)=d[\langle{\hat{x}}\rangle]/{dt}$ under OBCs where $[\langle{\hat{x}}\rangle]=2\sum_{l}l[\langle\psi_{t}|\hat{n}_{l}|\psi_{t}\rangle]/L$ is the average position of the particles [73]. If we assume that the velocity of a single left-moving particle is $v_{0}$, then a fraction of $2|v_{0}|t/L$ fermionic particles will reach the left side at time $t$ ($t<t_{c}$). As a result, only the remaining particles contribute to the velocity. The average velocity before the transition is thus given by ${v}(t)=\left(1-2|v_{0}|t/L\right)v_{0}$. We plot ${v}(t)$ in the inset of Fig. 1(b), where the black dashed line represents a linear fit of the data. The slope of the line is determined to be $0.8012v^{2}_{0}=$ 0.801, from which we obtain $v_{0}=-0.633$. Since the particles with the velocity $v_{0}$ are within the bulk, their properties (including velocities) should be determined by the system under PBCs as we have expected. We now use $v_{0}=({2}/{L})\sum_{k}{n}_{k}v_{k}$ to estimate $v_{0}$ through pure bulk dynamics. Here, $v_{k}=-2\sin{k}$ is the group velocity of the free fermion Hamiltonian at the momentum $k$, and ${n}_{k}=[\langle\psi_{\mathrm{steady}}|\hat{c}_{k}^{\dagger}\hat{c}_{k}|\psi_{\mathrm{steady}}\rangle]$ is the density of particles at $k$ in the steady state $|\psi_{\mathrm{steady}}\rangle$ of the dynamics under PBCs [see Fig. 2(c)]. The estimated velocity $v_{0}$ is $-0.633$, which is the same as the velocity calculated under OBCs. The critical rescaled time $t_{c}/L$ is determined by the fact that all particles move to the left side at this time so that $v(t_{c})=0$. Thus, $t_{c}/L=1/(2|v_{0}|)=0.79$, which is in excellent agreement with the results in Fig. 1(a). The transition is reminiscent of the kinetic phase transition where the average velocity vanishes with respect to a control parameter such as noise [75, 76]. In contrast, our transition occurs with respect to a rescaled time.
Dynamical transition with quasiperiodic potential.— We now investigate the dynamical entanglement transition that occurs in the presence of the quasiperiodic potential. By numerically analyzing the entanglement entropy $S_{L/2}$ at different times and for various values of $L$ and $W$, we map out the phase diagram in Fig. 3(a). We see that as $W$ increases, the dynamical entanglement transition persists, with the transition point being delayed. Specifically, we plot the evolution of $S_{L/2}$ for various values of $L$ when $W=1$ in Fig. 3(b), clearly illustrating that $S_{L/2}$ transitions from the logarithmic scaling [see the inset of Fig. 3(b)] to the area-law scaling. At the transition point, $t_{c}/L\approx 1.63$, which is larger than the transition point value in the case without quasidisorder. The delayed transition point is attributed to the fact that quasidisorder slows down the particles’ motion, but eventually, the particles migrate to the left boundary. In addition, we find that the probability of the evolving state being in an ideal many-body skin state suddenly rises across the transition point [73], similar to the case without disorder.
To further provide evidence for the dynamical transition, we calculate the connected density-density correlation function $C(l)=[\langle\hat{n}_{L/2}\rangle\langle\hat{n}_{L/2+l}\rangle-\langle\hat{n}_{L/2}\hat{n}_{L/2+l}\rangle]$ for different values of $W$ and $t/L$ and plot them in Fig. 3(c). We see clearly that $C(l)$ decays algebraically when the point $(W,t/L)$ falls within the log-law regime, whereas it decays exponentially in the area-law regime. This change in the behavior of $C(l)$ further reveals the existence of the dynamical transition.
Quasidisorder induced entanglement phase transition.— Previous studies have shown that disorder or quasidisorder can drive an entanglement phase transition in the steady state of the non-Hermitian Hamiltonian dynamics [34, 35, 36]. In our case, we find that the feedback measurement dynamics under OBCs always leads to a skin state for the steady state even for a large $W$ [see Fig. 3(d)], thereby excluding the existence of the phase transition in the steady state. Furthermore, Fig. 4(a) shows that when quasidisorder is sufficiently strong, $S_{L/2}$ directly approaches a steady value as time evolves, in contrast to the small $W$ case where the evolution exhibits a dynamical transition, as discussed earlier. In other words, the dynamical transition with respect to $t/L$ disappears for a large $W$.
Figure 4: (a) The time evolution of the trajectory averaged entanglement entropy $S_{L/2}$ for a system under OBCs with $L=64,128,192,256,320$ (from light to dark colors) at $W=0.6$ (blue lines) and $W=4$ (red lines). (b) The maximum entanglement entropy $S_{L/2}^{\mathrm{max}}$ and (c) the maximum mutual information $I_{AB}^{\mathrm{max}}$ between two subsystems $A$ and $B$ with respect to $W$ for different system sizes. In (b), the inset plots $S^{\mathrm{max}}_{L/2}$ versus $L$ in the linear-log scale at different $W$ with a log-law and area-law fit (the dashed lines), and in (c), the inset displays the positions and sizes of the two subsystems. The mutual information is defined as $I_{AB}=S_{A}+S_{B}-S_{A\cup B}$, where $S_{A}$ and $S_{B}$ are entanglement entropy of the subsystems A and B, respectively, and $S_{A\cup B}$ is the joint entanglement entropy. In (b) and (c), the maximum value refers to the peak of the corresponding trajectory averaged quantities during time evolution. (d) The correlation function $C(l)$ versus $(L/\pi)\sin{(\pi l/L)}$ at $W=0.8,4,6$ with $L=320$. The inset displays a data collapse for different system sizes $L=64,128,192,256,320$.
To quantitatively identify the quasidisorder induced disappearance of the dynamical transition, we calculate the maximum value $S_{L/2}^{\mathrm{max}}$ of the entanglement entropy as time evolves at each $W$ and plot $S_{L/2}^{\mathrm{max}}$ versus $W$ in Fig. 4(b). The figure shows that $S_{L/2}^{\mathrm{max}}$ first rises to a peak near $W\approx 0.8$ and then collapses to the same curve, suggesting the entanglement transition from a log-law scaling to an area-law scaling as $W$ increases [see the inset in Fig. 4(b)]. We find that the transition of $S_{L/2}^{\mathrm{max}}$ under OBCs reveals the entanglement transition in the steady state of the dynamics under PBCs [73]. It arises because for weak disorder, at short times before the transition to the area-law regime, the dynamics is mainly governed by the system under PBCs. For strong disorder, since both skin states and localized states respect the area-law entanglement scaling, the entanglement dynamics under OBCs shows a similar growth behavior as its dynamics under PBCs.
We further plot the maximum mutual information $I_{AB}^{\mathrm{max}}$ between two subsystems A and B [see the inset of Fig. 4(c)] in the dynamics under OBCs as a function of the quasidisorder strength $W$ in Fig. 4(c). Similar to the entanglement entropy, we see that $I_{AB}^{\mathrm{max}}$ first develops a peak near $W\approx 0.8$ and then declines to zero, providing additional evidence for the quasidisorder induced transition. We also observe that while $I_{AB}^{\mathrm{max}}$ declines to zero at a smaller $W$ for a larger system, the two lines representing the two largest systems ($L=448,480$) almost overlap, suggesting a possible smooth transition in the thermodynamic limit. To gain further insights into this transition, we also calculate the correlation function $C(l)$ of the evolving state at the time when $S_{L/2}$ reaches its maximum value. We find that for weak disorder, the correlation function exhibits a power-law decay, whereas for strong disorder, it shows an exponential decay [see Fig. 4(d)].
In summary, we have demonstrated a novel dynamical transition with respect to a rescaled time in a continuously monitored free fermion system under OBCs. We show that the transition arises from the competition between the bulk dynamics and boundary skin effects. Our findings are expected to have broader applicability beyond our specific scenario, as other systems described by the Lindblad master equation can also exhibit the skin effects [37, 38, 39]. In addition, we find that while the steady state is still the skin state in the presence of strong quasidisorder, the maximum entanglement exhibits a transition from a logarithmic to an area-law scaling as the quasidisorder strength increases. Since onsite disorder is ubiquitous, we have provided the results for the case with onsite disorder (rather than quasidisorder) in the Supplemental Material, showing that the dynamical transition persists. We note that the dynamical transition also occurs for other initial states [73]. Our work is also reminiscent of the phenomenon of a metastable state decaying to a steady state in Markovian open quantum dynamics [4]. However, the latter does not necessarily correspond to a sharp transition [73]. Finally, we propose a quantum circuit to observe the novel dynamical transition in the Supplemental Material [73], which can be realized with superconducting qubits [78], trapped ions [79] or Rydberg atoms [80, 81].
Acknowledgements.
We thank J.-H. Wang, X. Feng, S. Liu, and X. Gao for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 11974201), Tsinghua University Dushi Program and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University